### v1

@@ 4 Discussion
We derive the velocity dispersion functions of pressure-dominated galaxies in five EAGLE simulations with various AGN feedback models implemented. The shapes of VDFs measured from these simulations vary; particularly, the VDF measured from EAGLE-NoAGN shows a much flatter slope indicating the relative effective formation of high velocity dispersion galaxies. We also compare these simulated VDFs with the observed VDF. To understand the implications of the differences in VDF shapes, we first examine the impact of sample galaxy selection in Section 4.1. We then discuss the role of AGN feedback to interpret the implication of differences in simulated VDFs in Section 4.2.
4.1 Galaxy Sample Selection
@@
Figure 7 shows the specific star formation rate (sSFR) as a function of $v/\sigma$ for subhalos for $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$ in the five EAGLE simulations. In Figure 7, galaxies below the horizontal line are classified as quiescent galaxies, identified using the same selection criteria as in Sohn2024b (hereafter, the quiescent galaxy sample). Galaxies to the left of the vertical line are pressure-dominated systems (hereafter, the pressure-dominated sample). The overlap between the quiescent and pressure-dominated samples is surprisingly small. Among quiescent galaxies, $\sim 75-85\penalty 10000\ \%$ are pressure-dominated galaxies, whereas only $\sim 30-50\penalty 10000\ \%$ of pressure-dominated galaxies are quiescent galaxies. This difference in sample selection influences the shape of the resulting VDFs.
@@
Figure 8: The median $M_{*}-\sigma_{*}$ relation for EAGLE-50. Red circles and blue triangles show the relations derived from pressure-dominated galaxies with $v/\sigma<0.5$ and the quiescent galaxies with sSFR $2\times 10^{-11}$ yr-1, respectively. The error bars indicate the $1\sigma$ standard deviations.
@@
Figure 8 compares the $M_{*}-\sigma_{*}$ relations for the pressure-dominated sample ($v/\sigma<0.5$, red circles) and the quiescent sample (sSFR $2\times 10^{-11}$ yr-1, blue triangles) in EAGLE-50 simulation. The median $M_{*}-\sigma_{*}$ relations for the two samples are consistent over the mass range we explore. In other words, the stellar velocity dispersion does not change significantly at fixed stellar mass, regardless of how galaxies are identified. We also note that the pressure-dominated and the quiescent samples in other EAGLE simulations are also consistent without any offset.
@@
Figure 9: Comparison between VDFs constructed based on galaxy samples identified based on the $v/\sigma$ and the sSFR selections in five EAGLE simulations. Black circles show the observed VDF derived from SDSS field galaxies at $0.03\leq z\leq 0.10$ (Sohn2017b). Symbols are identical from Figure 8. Error bars on the solid lines indicate Poisson errors. We display only the upper values with arrows for bins including fewer than 10 galaxies.
@@
Finally, Figure 9 compares the VDFs derived from two different samples: the quiescent galaxy sample (blue triangles) and the pressure-dominated sample (red circles). The VDFs are generally consistent at $\log\sigma_{*}>2.0$, indicating that differences in galaxy classification have minimal impact on the VDF shape at high velocity dispersions. However, at $\log\sigma_{*}<2.0$, the VDF from the pressure-dominated sample exhibits a significantly steeper slope than that of the quiescent sample. This steepening is due to the inclusion of many pressure-dominated galaxies with ongoing star formation (i.e., galaxies in the upper left region of Figure 7). More importantly, our analysis indicates that differences in sample selection alone cannot account for the substantial discrepancy between the observed and simulated VDFs, regardless of the adopted AGN feedback model.
@@ 4.2 The Role of AGN Feedback in Velocity Dispersion
Our tests imply two important aspects of the VDFs derived from EAGLE simulations with various AGN feedback models. First, the VDFs derived from EAGLE simulations are significantly offset from the observed VDFs, similar to the VDFs derived from other simulations including TNG100 and TNG300 (Sohn2024b). Second, the $M_{*}-\sigma_{*}$ relation and the VDF vary little among EAGLE simulations, except for the one with no AGN feedback model. The steeper $M_{*}-\sigma_{*}$ relation of the galaxies in the EAGLE-NoAGN simulation is intriguing because the detailed physical mechanism of the AGN feedback affecting the stellar kinematics is unclear.
@@
To understand the role of AGN feedback on the stellar kinematics, we examine the origin of the peculiarities in the $M_{*}-\sigma_{*}$ relation and the VDF from EAGLE-NoAGN simulation. Unlike the $M_{*}-\sigma_{*}$ relation we show in Figure 4, we now investigate the the stellar mass to $\sigma_{*}$ relation, but based on the $M_{*}$ measured within 10 kpc aperture (i.e., $M_{*,z,10\penalty 10000\ \rm kpc}$).
@@
Figure 10: The median stellar velocity dispersions measured within a cylindrical volume within a 10 kpc aperture as a function of the stellar mass measured within a 10 kpc aperture for pressure-dominated galaxies in various EAGLE simulations at $z=0$. Blue triangles, red diamonds, and green solid line show the relations from EAGLE-50, EAGLE-eAGN-dT, and EAGLE-NoAGN, respectively. The error bar and shaded region indicate the $1\sigma$ standard deviation of the galaxies in each stellar mass bin.
@@
In Figure 10, the $M_{*,10\penalty 10000\ \rm kpc}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations derived from three simulations (i.e., EAGLE-50, EAGLE-eAGN-dT, and EAGLE-NoAGN) do not differ from each other, unlike the $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations shown in Figure 4. The discrepancy between the $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ and $M_{*,z,10\penalty 10000\ \rm kpc}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations indicates that 1) the stellar velocity dispersion is directly proportional to the stellar mass within the aperture we measure the velocity dispersion and 2) the stellar mass distribution within galaxies varies depending on the AGN feedback models (resulting in the different slopes in $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations). Figure 11 shows that the mass concentration of simulated galaxies varies depending on the adopted AGN feedback model. For $\log M_{*}<10.5$, the galaxies in the EAGLE-NoAGN (the green line) continue to follow a proportional relation with $M_{*}$, whereas galaxies in the other EAGLE simulations show systematically lower $\log M_{*,z,10\penalty 10000\ \rm kpc}$, reflecting reduced mass concentrations within 10 kpc due to AGN feedback.
@@
Figure 11: The median stellar mass measured within a cylindrical volume with a 10 kpc aperture as a function of the total stellar mass of pressure-dominated galaxies in various EAGLE simulations. Symbols are the same in Figure 10. The dashed line indicates the relations between total stellar mass and the stellar mass within half-mass radius. The dotted line indicates the one-to-one correspondence between the two stellar masses. The error bar and shaded region indicate the $1\sigma$ standard deviation of the galaxies in each stellar mass bin.
@@
Several physical mechanisms contribute to variations in the stellar mass distribution within galaxies (e.g., Furlong2017; Choi2018). First, galaxies naturally lose stellar mass as old stars through supernova explosions or during the asymptotic giant branch (AGB) phase. Second, stellar mass redistribution occurs through mergers or interactions between galaxies. These two processes likely operate to a similar extent regardless of the AGN feedback model employed.
@@
The AGN feedback can change the stellar mass distribution by disturbing the gravitational potential (e.g., Choi2018). When AGN feedback expels gas from galaxy centers, it reduces the central gravitational potential, causing the inner regions of galaxies to ‘puff up’. As a result, both the stellar and dark matter distributions expand, and the central potential becomes shallower (adiabatic expansion, e.g., Hopkins2010; Fan2008; Fan2010). This ‘puff up’ process generally increases the size of galaxies due to the migration of stars (Furlong2017; Choi2018). Galaxies in simulations with no AGN feedback then are compact compared to their counterparts in simulations with AGN feedback. Indeed, galaxies in EAGLE-NoAGN are generally more compact than those in other EAGLE simulations that include AGN feedback, at a fixed stellar mass. In the EAGLE-NoAGN simulation, galaxies with $M_{*}>10^{11}\penalty 10000\ M_{\odot}$ have median half-mass radii of 3–30 kpc, whereas galaxies with similar stellar masses in the other EAGLE simulations are much larger (10 – 130 kpc), with median sizes about four times greater.
@@
Figure 12 shows that the increase in stellar mass in the central region is strongly affected by AGN feedback. Figure 12 (a) shows the change of stellar mass within a 10 kpc aperture as a function of time. We plot the median variation of stellar mass of pressure-dominated galaxies at $z<1$ in five EAGLE simulations. In general, galaxies in all EAGLE simulations show continuous mass growth at <10<10 kpc. However, the median stellar mass growth rate varies significantly depending on the AGN feedback models. The mass growth rate of galaxies simulated with the AGN feedback is generally decreasing as a function of time. In contrast, the mass growth rate of galaxies in EAGLE-NoAGN exhibits a monotonic increase across the redshift range we explore.
@@
Figure 12: (a) The median change of $M_{*,z,10\penalty 10000\ \rm kpc}$ and (b) The median of $M_{*,z,10\penalty 10000\ \rm kpc}$ as a function of redshift in various EAGLE simulations at $z<1$. The symbols are the same in Figure 5.
@@
Figure 13: (a) The median sSFR within a spherical 10 kpc aperture in the five EAGLE simulations as a function of redshift. (b) The median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$, the amount of stellar mass within a 10 kpc spherical aperture increased by in-situ star formation, in various EAGLE simulations as a function of redshift at $z<1$. The symbols are the same in Figure 5.
@@
The most significant change in the stellar mass distribution due to the AGN feedback results from the variation in the star formation rate. Figure 13 (a) displays the median specific star formation rate (sSFR) of pressure-dominated galaxies in EAGLE simulations. We note that the specific star formation rate is measured within a spherical 10 kpc aperture for the galaxies, instead of the cylindrical aperture as we examined for Figure 12. The median specific star formation rate (sSFR) of galaxies in EAGLE-NoAGN is higher than that of galaxies in other EAGLE simulations. Overall, the sSFR of galaxies decreases over time, but the sSFR of galaxies in EAGLE-NoAGN decreases relatively slowly than for galaxies in other EAGLE simulations.
@@
Figure 13 (b) compares the amount of stellar mass within a 10 kpc spherical aperture increased by in-situ star formation (i.e., $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ kpc}$) at each redshift. To compute the in-situ stellar mass increment, we simply assume a linear variation of the star formation rate of galaxies between each redshift, then we compute the mass of newly formed stars in between the snapshots. The median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ of galaxies in EAGLE-NoAGN remains constant at $z<1$, while the median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ of galaxies in other EAGLE simulations continuously decrease over time. The decrease in $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ observed in most EAGLE simulations is clearly a result of the reduced availability of cool gas due to AGN feedback.
@@
Additionally, we examine the ex situ stellar mass growth within a 10 kpc aperture. In the four EAGLE simulations that include AGN feedback, the ex situ mass growth is typically an order of magnitude larger than the in situ mass growth. In contrast, galaxies in the EAGLE-NoAGN simulation show comparable in situ and ex situ mass growth at $z<1$. This indicates that the absence of AGN feedback leads to substantial stellar mass growth from internal star formation in the central region, comparable to that driven by mergers or accretion.
@@
Figure 12 and Figure 13 illustrate the different stellar mass evolution and evolution of stellar mass distribution of galaxies in EAGLE simulations with various AGN feedback models. In general, galaxies in all EAGLE simulations show stellar mass growth at $z<1$ in the central region. The growth rate varies depending on the implemented AGN feedback model; the galaxies in EAGLE-NoAGN show much significant mass growth in the central region. When the AGN feedback is implemented, the central black hole activity suppresses star formation in the central region by heating the gas and expulsion of gas. As a result, the stellar mass growth in the central region is suppressed, in turn, the central stellar velocity dispersions that are proportional to the mass within the central region do not increase significantly. The difference between $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations and the VDFs derived from various EAGLE simulations originates from the different star formation history in the central region.

### v2

@@ 4 Discussion
We derived the VDFs of pressure-dominated galaxies in five EAGLE simulations with various AGN feedback models implemented. The shapes of VDFs measured from these simulations vary; particularly, the VDF measured from EAGLE-NoAGN shows a much flatter slope, indicating the relative effective formation of high-velocity dispersion galaxies. We also compared these simulated VDFs with the observed VDF. To understand the implications of the differences in VDF shapes, we first examine the impact of sample-galaxy selection in Section 4.1. We then discuss the role of AGN feedback in our interpretation of the implication of differences in simulated VDFs in Section 4.2.
4.1 Galaxy-sample selection
@@
Figure 7 shows the specific star formation rate (sSFR) as a function of $v/\sigma$ for subhalos for $M_{*}>10^{9}\penalty 10000\ \rm{M}_{\odot}$ in the five EAGLE simulations. In Figure 7, galaxies below the horizontal line are classified as quiescent galaxies, which were identified using the same selection criteria as in Sohn et al. (2024a) (hereafter the quiescent galaxy sample). Galaxies to the left of the vertical line are pressure-dominated systems (hereafter the pressure-dominated sample). The overlap between the quiescent and pressure-dominated samples is surprisingly small. Among quiescent galaxies, $\sim 75-85\penalty 10000\ \%$ are pressure-dominated galaxies, whereas only $\sim 30-50\penalty 10000\ \%$ of pressure-dominated galaxies are quiescent galaxies. This difference in sample selection influences the shape of the resulting VDFs.
@@
Figure 8: Median $M_{*}-\sigma_{*}$ relation for EAGLE-50. Red circles and blue triangles show the relations derived from pressure-dominated galaxies with $v/\sigma<0.5$ and the quiescent galaxies with sSFR $2\times 10^{-11}$ yr-1, respectively. The error bars indicate the $1\sigma$ standard deviations.
@@
Figure 8 compares the $M_{*}-\sigma_{*}$ relations for the pressure-dominated sample ($v/\sigma<0.5$, red circles) and the quiescent sample (sSFR $2\times 10^{-11}$ yr-1, blue triangles) in the EAGLE-50 simulation. The median $M_{*}-\sigma_{*}$ relations for the two samples are consistent over the mass range we explored. In other words, the stellar velocity dispersion does not change significantly at fixed stellar mass, regardless of how galaxies are identified. We also note that the pressure-dominated and quiescent samples in other EAGLE simulations are also consistent without any offset.
@@
Figure 9: Comparison between VDFs constructed from galaxy samples identified using $v/\sigma$-based and sSFR-based selections in five EAGLE simulations. Black circles show the observed VDF derived from SDSS field galaxies at $0.03\leq z\leq 0.10$ (Sohn et al., 2017b). Symbols are identical to those of Figure 8. Error bars on the solid lines indicate Poisson errors. We only display the upper values with arrows for bins including fewer than ten galaxies.
@@
Finally, Figure 9 compares the VDFs derived from two different samples: the quiescent galaxy sample (blue triangles) and the pressure-dominated sample (red circles). The VDFs are generally consistent at $\log\sigma_{*}>2.0$, indicating that differences in galaxy classification have minimal impact on the VDF shape at high-velocity dispersions. However, at $\log\sigma_{*}<2.0$, the VDF from the pressure-dominated sample exhibits a significantly steeper slope than that of the quiescent sample. This steepening is due to the inclusion of many pressure-dominated galaxies with ongoing star formation (i.e., galaxies in the upper left region of Figure 7). More importantly, our analysis indicates that differences in sample selection alone cannot account for the substantial discrepancy between the observed and simulated VDFs, regardless of the AGN feedback model adopted.
@@ 4.2 The role of AGN feedback in velocity dispersion
Our tests imply two important aspects of the VDFs derived from EAGLE simulations with various AGN feedback models. First, the VDFs derived from EAGLE simulations are significantly offset from the observed VDFs, similarly to the VDFs derived from other simulations including TNG100 and TNG300 (Sohn et al., 2024a). Second, the $M_{*}-\sigma_{*}$ relation and the VDF vary little among EAGLE simulations, except for the one with no AGN feedback model. The steeper $M_{*}-\sigma_{*}$ relation of the galaxies in the EAGLE-NoAGN simulation is intriguing because the detailed physical mechanism of the AGN feedback affecting the stellar kinematics is unclear.
@@
To understand the role of AGN feedback on the stellar kinematics, we examined the origin of the peculiarities in the $M_{*}-\sigma_{*}$ relation and the VDF from the EAGLE-NoAGN simulation. Unlike the $M_{*}-\sigma_{*}$ relation we show in Figure 4, we investigated the stellar mass-to-$\sigma_{*}$ relation, but based on the $M_{*}$ measured within a 10 kpc aperture (i.e., $M_{*,z,10\penalty 10000\ \rm kpc}$).
@@
Figure 10: Median stellar velocity dispersions measured within a cylindrical volume within a 10 kpc aperture as a function of the stellar mass measured within a 10 kpc aperture for pressure-dominated galaxies in various EAGLE simulations at $z=0$. Blue triangles, red diamonds, and the solid green line show the relations from EAGLE-50, EAGLE-eAGN-dT, and EAGLE-NoAGN, respectively. The error bar and shaded region indicate the $1\sigma$ standard deviation of the galaxies in each stellar mass bin.
@@
In Figure 10, the $M_{*,10\penalty 10000\ \rm kpc}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations derived from three simulations (i.e., EAGLE-50, EAGLE-eAGN-dT, and EAGLE-NoAGN) do not differ from each other, unlike the $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations shown in Figure 4. The discrepancy between the $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ and $M_{*,z,10\penalty 10000\ \rm kpc}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations indicates that 1) the stellar velocity dispersion is directly proportional to the stellar mass within the aperture where we measured the velocity dispersion; and 2) the stellar mass distribution within galaxies varies depending on the AGN feedback models (resulting in the different slopes in $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations). Figure 11 shows that the mass concentration of simulated galaxies varies depending on the adopted AGN feedback model. For $\log M_{*}<10.5$, the galaxies in the EAGLE-NoAGN (the green line) continue to follow a proportional relation with $M_{*}$, whereas galaxies in the other EAGLE simulations show systematically lower $\log M_{*,z,10\penalty 10000\ \rm kpc}$, reflecting reduced mass concentrations within 10 kpc due to AGN feedback.
@@
Figure 11: Median stellar mass measured within cylindrical volume with 10 kpc aperture as function of total stellar mass of pressure-dominated galaxies in various EAGLE simulations. Symbols are the same as in Figure 10. The dashed line indicates the relations between total stellar mass and the stellar mass within half-mass radius. The dotted line indicates the one-to-one correspondence between the two stellar masses. The error bar and shaded region indicate the $1\sigma$ standard deviation of the galaxies in each stellar mass bin.
@@
Several physical mechanisms contribute to variations in the stellar mass distribution within galaxies (e.g., Furlong et al., 2017; Choi et al., 2018). First, galaxies naturally lose stellar mass as old stars through supernova explosions or during the asymptotic-giant-branch (AGB) phase. Second, stellar mass redistribution occurs through mergers or interactions between galaxies. These two processes likely operate to a similar extent regardless of the AGN feedback model employed.
@@
The AGN feedback can change the stellar mass distribution by disturbing the gravitational potential (e.g., Choi et al., 2018). When AGN feedback expels gas from galaxy centers, it reduces the central gravitational potential, causing the inner regions of galaxies to ”puff up.” As a result, both the stellar and DM distributions expand, and the central potential becomes shallower (adiabatic expansion; e.g., Hopkins and Quataert, 2010; Fan et al., 2008, 2010). This puff-up process generally increases the size of galaxies due to the migration of stars (Furlong et al., 2017; Choi et al., 2018). Galaxies in simulations with no AGN feedback are then compact compared to their counterparts in simulations with AGN feedback. Indeed, galaxies in EAGLE-NoAGN are generally more compact than those in other EAGLE simulations that include AGN feedback, at a fixed stellar mass. In the EAGLE-NoAGN simulation, galaxies with $M_{*}>10^{11}\penalty 10000\ M_{\odot}$ have median half-mass radii of 3 – 30 kpc, whereas galaxies with similar stellar masses in the other EAGLE simulations are much larger (10 – 130 kpc), with median sizes about four times greater.
@@
Figure 12 shows that the increase in stellar mass in the central region is strongly affected by AGN feedback. Figure 12 (a) shows the change of stellar mass within a 10 kpc aperture as a function of time. We plot the median variation of stellar mass of pressure-dominated galaxies at $z<1$ in five EAGLE simulations. In general, galaxies in all EAGLE simulations show continuous mass growth at <10<10 kpc. However, the median stellar mass growth rate varies significantly depending on the AGN feedback models. The mass-growth rate of galaxies simulated with the AGN feedback generally decreases as a function of time. In contrast, the mass-growth rate of galaxies in EAGLE-NoAGN exhibits a monotonic increase across the redshift range we explored.
@@
Figure 12: (a) Median change of $M_{*,z,10\penalty 10000\ \rm kpc}$ and (b) median of $M_{*,z,10\penalty 10000\ \rm kpc}$ as a function of redshift in various EAGLE simulations at $z<1$. The symbols are the same as in Figure 5.
@@
Figure 13: (a) Median sSFR within spherical 10 kpc aperture in five EAGLE simulations as a function of redshift. (b) The median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$, the amount of stellar mass within a 10 kpc spherical aperture increased by in situ star formation in various EAGLE simulations as a function of redshift at $z<1$. The symbols are the same as in Figure 5.
@@
The most significant change in the stellar mass distribution due to the AGN feedback results from the variation in the star formation rate. Figure 13 (a) displays the median sSFR of pressure-dominated galaxies in EAGLE simulations. We note that the sSFR is measured within a spherical 10 kpc aperture for the galaxies instead of the cylindrical aperture as we examined for Figure 12. The median sSFR of galaxies in EAGLE-NoAGN is higher than that of galaxies in other EAGLE simulations. Overall, the sSFR of galaxies decreases over time, but the sSFR of galaxies in EAGLE-NoAGN decreases relatively slowly compared to that of galaxies in other EAGLE simulations.
@@
Figure 13 (b) compares the amount of stellar mass within a 10 kpc spherical aperture increased by in situ star formation (i.e., $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ kpc}$) at each redshift. To compute the in situ stellar mass increment, we simply assumed a linear variation of the star formation rate of galaxies between each redshift; then we computed the mass of newly formed stars in between the snapshots. The median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ of galaxies in EAGLE-NoAGN remains constant at $z<1$, while the median $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ of galaxies in other EAGLE simulations continuously decrease over time. The decrease in $M_{*,\penalty 10000\ \rm SF,10\penalty 10000\ \rm kpc}$ observed in most EAGLE simulations is clearly a result of the reduced availability of cool gas due to AGN feedback.
@@
Additionally, we examined the ex situ stellar mass growth within a 10 kpc aperture. In the four EAGLE simulations that include AGN feedback, the ex situ mass growth is typically an order of magnitude larger than the in situ mass growth. In contrast, galaxies in the EAGLE-NoAGN simulation show comparable in situ and ex situ mass growth at $z<1$. This indicates that the absence of AGN feedback leads to substantial stellar mass growth from internal star formation in the central region that is comparable to that driven by mergers or accretion.
@@
Figure 12 and Figure 13 illustrate the different stellar mass evolution and evolution of stellar mass distribution of galaxies in EAGLE simulations with various AGN feedback models. In general, galaxies in all EAGLE simulations show stellar mass growth at $z<1$ in the central region. The growth rate varies depending on the implemented AGN feedback model; the galaxies in EAGLE-NoAGN show much significant mass growth in the central region. When the AGN feedback is implemented, the central black hole activity suppresses star formation in the central region by heating the gas and expulsion of gas. As a result, the stellar mass growth in the central region is suppressed; in turn, the central stellar velocity dispersions that are proportional to the mass within the central region do not increase significantly. The difference between $M_{*}-\sigma_{*,z,10\penalty 10000\ \rm kpc}$ relations and the VDFs derived from various EAGLE simulations originates from the different star formation history in the central region.